OpenStudy (anonymous):

Prove: sin θ - sin θ•cos2 θ = sin3 θ. You must show all work.

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

Is cos2theta ''cos to the power 2 theta'' or ''cos of 2theta''?

OpenStudy (anonymous):

i feel something is missing in it.

OpenStudy (anonymous):

thats the exact question so thats all they gave me

OpenStudy (michele_laino):

we have to apply this identity, at the left side: \[\large {\left( {\cos \theta } \right)^2} = 1 - {\left( {\sin \theta } \right)^2}\]

OpenStudy (michele_laino):

so, you should get this: \[\large \sin \theta - \sin \theta {\left( {\cos \theta } \right)^2} = \sin \theta - \sin \theta \left[ {1 - {{\left( {\sin \theta } \right)}^2}} \right] = ...?\]

OpenStudy (anonymous):

\[\sin(\theta)-\sin(\theta)(\cos^2(\theta)=\sin(\theta)[1-\cos^2(\theta)]=\sin(\theta)(\sin^2(\theta)=\sin^3(\theta)\]

OpenStudy (anonymous):

sin^3(0)

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

it should be sin^3(theta) else the equation become 0=sin^3(theta)

OpenStudy (anonymous):

so that the answer??

OpenStudy (anonymous):

obliously sin^3(theta)

OpenStudy (michele_laino):

yes! Since we have showed, that the left side is equal to the right side

OpenStudy (anonymous):

thankyou guys

OpenStudy (michele_laino):

thank you!

OpenStudy (anonymous):

you are welcome.!